What is the magnitude of the electric field produced by a charge of magnitude 5.20 uC at a distance of (a) 1.00 m and (b) 3.00 m? N/C Submit Part B N/C Submit
Answer #1

Solution:

(a) 1.00 m

Answer: The magnitude of the electric field is

E=kq/r^{2}

Where:

q = 5.20 × 10-6C

.. E = 9 × 109 × 5.20 × 10-6/(1.00m)

  4.68 × 104 N/C

(b) 3.00 m

Answer: The magnitude of the electric field is

E=kq/r^{2}

Where:

q = 5.20 × 10-6C

.. E = 9 × 10 × 5.20 × 10-6/(3.00m)

  5.20 x 103 N/C